# Generalizing Goodstein’s theorem and Cichon’s independence proof

### v1

@@ 3 Goodstein process
Now that Theorem 2.14 is established, Cichon’s trick applies as layed out in the introduction, and we obtain the reformulation of the generalized Goodstein principle (3). Since the function defined in (4) is expressed in terms of the Hardy hierarchy as in (6), if the generalized Goodstein principle were provable in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, we would obtain a proof of the totality of $H_{\alpha}$ for $\alpha:={\operatorname{T}}\cap\Omega_{1}$ in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using the canonical fundamental sequence $\alpha[n]:=\vartheta_{0}(\ldots\vartheta_{n}(0)\ldots)$, which is completely contained in each ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$}, so that $H_{\alpha}(k)=H_{\alpha[k]}(k)$, despite the well-known fact that $H_{\alpha}$ is not provably total in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$. We thus obtain
@@
Any ordinal $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ is maximal among all $\beta\in{\operatorname{T}}\cap\Omega_{1}$ such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. This is seen as follows: assuming otherwise let $\beta>\alpha$ be minimal such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. Then $\beta[k]<\alpha<\beta$, so that according to Lemma 2.13 it would follow that ${\mathrm{imc}}(\alpha)\geq k$, which is not the case as $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$.

### v2

@@ 3 Goodstein process
Now that Theorem 2.21 is established, Cichon’s trick applies as layed out in the introduction, and we obtain the reformulation of the generalized Goodstein principle (3). Since the function defined in (4) is expressed in terms of Hardy hierarchy as in (6), if the generalized Goodstein principle were provable in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, we would obtain a proof of the totality of $H_{\alpha}$ for $\alpha:={\operatorname{T}}\cap\Omega_{1}$ in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using the canonical fundamental sequence $\alpha[n]:=\vartheta_{0}(\ldots\vartheta_{n}(0)\ldots)$, which is completely contained in each ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$}, so that $H_{\alpha}(k)=H_{\alpha[k]}(k)$, despite the well-known fact that $H_{\alpha}$ is not provably total in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$. We thus obtain
@@
Any ordinal $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ is maximal among all $\beta\in{\operatorname{T}}\cap\Omega_{1}$ such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. This is seen as follows: assuming otherwise, let $\beta>\alpha$ be minimal such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. Then $\beta[k-1]<\alpha<\beta$, so that according to Lemma 2.17 it would follow that ${\mathrm{imc}}(\alpha)\geq k$, which is not the case as $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$.